(-5,3),(-2,6) in Ax+By=c form
Answer by Ave(106)

first find the slope
(6-3)/-2-(-5)
3/3
1
now use point slope form with m=1, (-2, 6)
y-6=1(x-(-2))
y-6=x+2 add a 6 and -x to both sides
-x+y=8 or
x-y=-8